In a survey of 468 registered voters, 152 of them wished to see Mayor Waffleskate lose her next election. The Waffleskate campaign claims that no more than 32% of registered voters wish to see her defeated. Does the 95% confidence interval for the proportion support this claim? (Hint: you should first construct the 95% confidence interval for the proportion of registered voters who whish to see Waffleskate defeated.)
Answer by Boreal(15235)

95% CI half interval for this uses p hat=152/468=0.3248
half-interval is 1.96*sqrt(0.3248*0.6752/468)=0.0424
the full interval to 3 places is (0.282, 0.367), and since 32% is in this interval, the claim cannot be disproven by the CI.